# On $\mathcal{F}$-semi-transitive families of operators

### v5

@@ Abstract.
Motivated by the concept of disjoint topological transitivity for supercyclicity given in [31], in this paper, we present the concept of disjoint Furstenberg-semi-transitivity. Then we provide sufficient conditions for cosine operator functions generated by adjoints of weighted composition operators to be Furstenberg-semi-transitive on the space of Radon measures on a locally compact, non-compact Hausdorff space. Also, we characterize disjoint Furstenberg-semi-transitive compositions of a right multiplier and a generalized bilateral weighted shift on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space. As a special case, we consider the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of the set of natural numbers. Finally, we illustrate our results with concrete examples.
@@ 1. Introduction
Now, the research on the dynamics of weighted composition operators has started from pioneering papers by Salas on the dynamics on weighted shifts, see [35, 36]. Since then, numerous papers on the dynamics of weighted composition operators on various function spaces have been published. In particular, the dynamics of cosine operator functions generated by weighted composition operators has been studied in for instance [10, 11, 12, 13, 14, 24, 38]. Moreover, the dynamics of cosine operator functions generated by some other operators has been investigated in for instance [23, 28, 29]. Finally, quite recently, topological transitivity of the adjoints of weighted composition operators has been investigated and studied in [23]. These adjoints considered in [23] act on the space of Radon measures on a locally compact, non-compact Hausdorff space. Although there are several published papers on the dynamics of cosine operator functions, supercyclicity of cosine operator functions has not been studied in any of these published papers. All these facts naturally lead us to study in this paper supercyclicity and $\mathcal{F}$-semi-transitivity (as a generalization of supercyclicity) of cosine operator functions generated by the adjoints of weighted composition operators. In Section 3 in this paper, after presenting the concept of $\mathcal{F}$-semi-transitivity, we obtain sufficient conditions for cosine operator functions generated by the adjoints of weighted composition operators to be $\mathcal{F}$-semi-transitive on the space of Radon measures (on a locally compact, non-compact Hausdorff space) equipped with the total variation norm (Theorem 3.1). A special case of this theory is the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of $\mathbb{N},$ as given in Corollary 3.2. We provide also concrete examples where the sufficient conditions of Theorem 3.1 and Corollary 3.2 in this paper are satisfied, but where the sufficient conditions of [23, Proposition 3.3] are not satisfied, illustrating thus the difference between (ordinary) topological transitivity and $\mathcal{F}$-semi-transitivity of cosine operator functions.
@@ 2. $(ii)$
Let $U_{1}$ and $U_{2}$ be two open non-empty subsets of $M(\Omega)$. Then $U_{1}\setminus\{0\}$ and $U_{2}\setminus\{0\}$ are also open and non-empty subsets of $M(\Omega)$. As in the proof of [23, Proposition 3.1] part $(ii)\Rightarrow(i)$, we can find two measures $\mu\in U_{1}\setminus\{0\}$, $\nu\in U_{2}\setminus\{0\}$, and a compact subset $K$ of $\Omega$ such that $|\mu|(K^{c})=|\nu|(K^{c})=0$. Also, we can find some $\delta>0$ such that $B(\mu,\delta)\subset U_{1}\setminus\{0\}$ and $B(\nu,\delta)\subset U_{2}\setminus\{0\}$. Choose $F\in\mathcal{F}$ and collections of Borel subsets $\{A_{z}\}_{z\in F}$, $\{D_{z}\}_{z\in F}$$\{E_{z}\}_{z\in F}$ satisfying the assumptions of $(ii)$ with respect to $\mu$, $\nu$, and
@@
for all $z\in F$. Now, by exactly the same arguments as in the proof of [23, Proposition 3.1] part $(ii)\Rightarrow(i)$, one can show that for each $z\in F$ we have
@@
hence the conditions of [23, Proposition 3.3] are not satisfied. This illustrates the difference between the sufficient conditions of Corollary 3.2 and the sufficient conditions of [23, Proposition 3.3].
@@ 4. Disjoint $\mathcal{F}$-semi-transitive generalized weighted shifts
Let $\mathcal{S}$ be a set, $\mathcal{F}$ be a family of subsets of $\mathcal{S}.$ and $\{T_{t,1}\}_{t\in S},\dots,\\ \{T_{t,N}\}_{t\in S}$ be families of bounded linear operators on a Banach space $X$. We say that $\{T_{t,1}\}_{t\in S},\dots,\{T_{t,N}\}_{t\in S}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ there exists some $\lambda_{t}\in\mathbb{R}^{+}$ satisfying that
@@
In this section, for a separable Hilbert space $H$ we let $B(H)$ be the space of all bounded linear operators on $H$, and $\mathcal{C}$ be the $C^{*}$-algebra of compact operators on $H$. For an orthonormal basis $\{e_{j}\}_{j\in\mathbb{Z}}$ for $H$, we let for each $m\in\mathbb{N}$, $P_{m}$ be the orthogonal projection onto $\text{Span}\{e_{-m},\dots,e_{m}\}$. We will denote by $\ell_{2}(\mathcal{C})$ the standard (right) Hilbert module over $\mathcal{C}$, see [32, Example 1.3.5]. Notice that $\ell_{2}(\mathcal{C})$ is a Banach algebra. Indeed, we can define multiplication on $\ell_{2}(\mathcal{C})$ as pointwise multiplication, i.e., if $\{x_{j}\}_{j\in\mathbb{Z}},\{y_{j}\}_{j\in\mathbb{Z}}\in\ell_{2}(\mathcal{C})$, then
@@
As shown in the proof of [26, Proposition 3.1], the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$.
@@
(2) For every $J,m\in\mathbb{N}$ and $\varepsilon>0$, there exists some $F\in\mathcal{F}$ and a sequence $\{S_{n}\}_{n\in\mathbb{N}}$ of bounded invertible linear operators on $H$ such that for each $j\in\mathbb{Z}$ with $J\leq j\leq J$ we can find families $\{D_{n,j}\}_{n\in F},\quad\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{G_{n,j}^{(N)}\}_{n\in F}$ of compact operators on $H$ satisfying for all $n\in F$ and $\ell,s\in\{1,\dots,N\}$ that
@@
The main idea in this proof is motivated by the proofs of [26, Theorem 3.2] and [27, Proposition 2.7]. Given $m\in\mathbb{N}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, for every $\eta\in\{-m,\dots,m\}$ we can find sequences
@@
By some calculations and arguments similar to those in the proof of [26, Theorem 3.2] and [27, Proposition 2.7], we can conclude that the sequences of operators
@@
Let $N=1,r_{1}=1$ and $H=L^{2}(\mathbb{R})$. As in [26, Example 3.3], choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$. Let $\{c_{j}\}_{j\in\mathbb{Z}}\subseteq\mathbb{R}^{+}$ be such that $c_{j}\geq C$ for all $j\in\mathbb{Z}$ and some $C>0$. For each $j\in\mathbb{Z}$ let $\alpha_{j}$ be the translation on $\mathbb{R}$ given by $\alpha_{j}(t)=t-c_{j}$ for all $t\in\mathbb{R}$. If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; then by the same arguments as in [26, Example 3.3] one can show that the conditions of Proposition 4.6 are satisfied when $W_{j}(t)=w_{j}\cdot(f\circ\alpha_{j})$ for all $f\in L^{2}(\mathbb{R}),j\in\mathbb{Z}$ and $H_{0}^{(j)}=H_{1}^{(j)}=C_{c}(\mathbb{R})$ for all $j\in\mathbb{Z}.$ However, by the construction of the sequence $\{w_{j}\}$, it follows that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}.$ On the other hand, if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then again the conditions of Proposition 4.6 will be satisfied, however, in this case we would again have by the construction of the sequence $\{w_{j}\}$ that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}.$
@@
for all $\{x_{i}\}_{i\in\mathbb{Z}}\in\ell_{2}(\mathcal{C}).$ By [32, Proposition 1.2.4 (i)], the operator $R_{s,n}$ is a bounded linear operator on $\ell_{2}(\mathcal{C})$ for each $s\in\{1,\dots,N\}$ and $n\in\mathbb{N}.$ Let $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ be the families of operators on $\ell_{2}(\mathcal{C})$ as defined in the beginning of this section. For each $s\in\{1,\dots,N\}$ and $n\in\mathbb{N},$ we let $T_{n,s}\circ R_{s,n}$ be the composition of $T_{n,s}$ and $R_{s,n},$ that is
@@
Let $\mathcal{O},V_{1},\dots,V_{N}$ be non-empty open subsets of $\ell_{2}(\mathcal{C})$. Then there exist some $J,m_{1}\in\mathbb{N}$ and some $x\in\mathcal{O}$, $y^{(1)}\in V_{1},\dots,y^{(N)}\in V_{N}$ such that $\tilde{p}_{J,m_{1}}x=x$ and $\tilde{p}_{J,m_{1}}y^{(\ell)}=y^{(\ell)}$ for each $\ell\in\{1,\dots,N\}.$ Since $\{P_{m}\}_{m\in\mathbb{N}}$ forms also a right approximate unit for $\mathcal{C}$ by [32, Proposition 2.2.1], it follows that we can find some $m_{2}\in\mathbb{N}$ such that $\tilde{p}_{J,m_{1}}x\tilde{p}_{J,m_{2}}\in\mathcal{O}$ and $\tilde{p}_{J,m_{1}}y^{(\ell)}\tilde{p}_{J,m_{2}}\in V_{l}$ for each $\ell\in\{1,\dots,N\}.$ Set $m:=\text{max}\{m_{1},m_{2}\}.$ Then $\tilde{p}_{J,m}x\tilde{p}_{J,m}\in\mathcal{O}$ and $\tilde{p}_{J,m}y^{(\ell)}\tilde{p}_{J,m}\in V_{l}$ for each $\ell\in\{1,\dots,N\}.$ Put $C_{y}=\sum_{s=1}^{N}\|y^{(s)}\|$ and

### v6

@@ Abstract.
Motivated by the concept of disjoint topological transitivity for supercyclicity given in [32], in this paper, we present the concept of disjoint Furstenberg-semi-transitivity. Then we provide sufficient conditions for cosine operator functions generated by adjoints of weighted composition operators to be Furstenberg-semi-transitive on the space of Radon measures on a locally compact, non-compact Hausdorff space. Also, we characterize disjoint Furstenberg-semi-transitive compositions of a right multiplier and a generalized bilateral weighted shift on the standard Hilbert module over the C*-algebra of compact operators on a separable Hilbert space. As a special case, we consider the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of the set of natural numbers. Finally, we illustrate our results with concrete examples. We end this paper with an algebraic generalization to the case of non-unital Banach algebras.
@@ 1. Introduction
Now, the research on the dynamics of weighted composition operators has started from pioneering papers by Salas on the dynamics on weighted shifts, see [36, 37]. Since then, numerous papers on the dynamics of weighted composition operators on various function spaces have been published. In particular, the dynamics of cosine operator functions generated by weighted composition operators has been studied in for instance [10, 11, 12, 13, 14, 25, 39]. Moreover, the dynamics of cosine operator functions generated by some other operators has been investigated in for instance [24, 29, 30]. Finally, quite recently, topological transitivity of the adjoints of weighted composition operators has been investigated and studied in [24]. These adjoints considered in [24] act on the space of Radon measures on a locally compact, non-compact Hausdorff space. Although there are several published papers on the dynamics of cosine operator functions, supercyclicity of cosine operator functions has not been studied in any of these published papers. All these facts naturally lead us to study in this paper supercyclicity and $\mathcal{F}$-semi-transitivity (as a generalization of supercyclicity) of cosine operator functions generated by the adjoints of weighted composition operators. In Section 3 in this paper, after presenting the concept of $\mathcal{F}$-semi-transitivity, we obtain sufficient conditions for cosine operator functions generated by the adjoints of weighted composition operators to be $\mathcal{F}$-semi-transitive on the space of Radon measures (on a locally compact, non-compact Hausdorff space) equipped with the total variation norm (Theorem 3.1). A special case of this theory is the case when the corresponding Furstenberg family under consideration is the family of all infinite subsets of $\mathbb{N},$ as given in Corollary 3.2. We provide also concrete examples where the sufficient conditions of Theorem 3.1 and Corollary 3.2 in this paper are satisfied, but where the sufficient conditions of [24, Proposition 3.3] are not satisfied, illustrating thus the difference between (ordinary) topological transitivity and $\mathcal{F}$-semi-transitivity of cosine operator functions.
@@
Section 5 provides algebraic generalization of Theorem 4.2 in this paper to the case of more general non-unital Banach algebras.
@@ 2. $(ii)$
Let $U_{1}$ and $U_{2}$ be two open non-empty subsets of $M(\Omega)$. Then $U_{1}\setminus\{0\}$ and $U_{2}\setminus\{0\}$ are also open and non-empty subsets of $M(\Omega)$. As in the proof of [24, Proposition 3.1] part $(ii)\Rightarrow(i)$, we can find two measures $\mu\in U_{1}\setminus\{0\}$, $\nu\in U_{2}\setminus\{0\}$, and a compact subset $K$ of $\Omega$ such that $|\mu|(K^{c})=|\nu|(K^{c})=0$. Also, we can find some $\delta>0$ such that $B(\mu,\delta)\subset U_{1}\setminus\{0\}$ and $B(\nu,\delta)\subset U_{2}\setminus\{0\}$. Choose $F\in\mathcal{F}$ and collections of Borel subsets $\{A_{z}\}_{z\in F}$, $\{D_{z}\}_{z\in F}$$\{E_{z}\}_{z\in F}$ satisfying the assumptions of $(ii)$ with respect to $\mu$, $\nu$, and
@@
for all $z\in F$. Now, by exactly the same arguments as in the proof of [24, Proposition 3.1] part $(ii)\Rightarrow(i)$, one can show that for each $z\in F$ we have
@@
hence the conditions of [24, Proposition 3.3] are not satisfied. This illustrates the difference between the sufficient conditions of Corollary 3.2 and the sufficient conditions of [24, Proposition 3.3].
@@ 4. Disjoint $\mathcal{F}$-semi-transitive generalized weighted shifts
Let $\mathcal{S}$ be a set, $\mathcal{F}$ be a family of subsets of $\mathcal{S}.$ and $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\\ \{T_{t,N}\}_{t\in\mathcal{S}}$ be families of bounded linear operators on a Banach space $X$. We say that $\{T_{t,1}\}_{t\in\mathcal{S}},\dots,\{T_{t,N}\}_{t\in\mathcal{S}}$ are disjoint $\mathcal{F}$-semi-transitive, or shortly $d\mathcal{F}$-semi-transitive, if for every collection of non-empty open subsets $\mathcal{O},V_{1},\dots,V_{N}$ of $X$, there exists some $F\in\mathcal{F}$ such that for all $t\in F$ there exists some $\lambda_{t}\in\mathbb{R}^{+}$ satisfying that
@@
In this section, for a separable Hilbert space $H$ we let $B(H)$ be the space of all bounded linear operators on $H$, and $\mathcal{C}$ be the $C^{*}$-algebra of compact operators on $H$. For an orthonormal basis $\{e_{j}\}_{j\in\mathbb{Z}}$ for $H$, we let for each $m\in\mathbb{N}$, $P_{m}$ be the orthogonal projection onto $\text{Span}\{e_{-m},\dots,e_{m}\}$. We will denote by $\ell_{2}(\mathcal{C})$ the standard (right) Hilbert module over $\mathcal{C}$, see [33, Example 1.3.5]. Notice that $\ell_{2}(\mathcal{C})$ is a Banach algebra. Indeed, we can define multiplication on $\ell_{2}(\mathcal{C})$ as pointwise multiplication, i.e., if $\{x_{j}\}_{j\in\mathbb{Z}},\{y_{j}\}_{j\in\mathbb{Z}}\in\ell_{2}(\mathcal{C})$, then
@@
As shown in the proof of [27, Proposition 3.1], the set $\{\tilde{p}_{J,m}\}_{(J,m)\in\mathbb{N}^{2}}$ forms a left approximate unit for $\ell_{2}(\mathcal{C})$.
@@
(2) For every $J,m\in\mathbb{N}$ and $\varepsilon>0$, there exists some $F\in\mathcal{F}$ and a sequence $\{S_{n}\}_{n\in F}$ of bounded invertible linear operators on $H$ such that for each $j\in\mathbb{Z}$ with $J\leq j\leq J$ we can find families $\{D_{n,j}\}_{n\in F},\quad\{G_{n,j}^{(1)}\}_{n\in F},\dots,\{G_{n,j}^{(N)}\}_{n\in F}$ of compact operators on $H$ satisfying for all $n\in F$ and $\ell,s\in\{1,\dots,N\}$ that
@@
The main idea in this proof is motivated by the proofs of [27, Theorem 3.2] and [28, Proposition 2.7]. Given $m\in\mathbb{N}$ and $j\in\mathbb{Z}$ with $J\leq j\leq J$, for every $\eta\in\{-m,\dots,m\}$ we can find sequences
@@
By some calculations and arguments similar to those in the proof of [27, Theorem 3.2] and [28, Proposition 2.7], we can conclude that the sequences of operators
@@
Let $N=1,r_{1}=1$ and $H=L^{2}(\mathbb{R})$. As in [27, Example 3.3], choose some $\{w_{j}\}_{j\in\mathbb{Z}}\subseteq L^{\infty}(\mathbb{R})$ such that $w_{j}>0$, $w_{j}^{-1}\in L^{\infty}(\mathbb{R})$ and $\parallel w_{j}\parallel\leq M$ for all $j\in\mathbb{Z}$ and some $M>1$. Let $\{c_{j}\}_{j\in\mathbb{Z}}\subseteq\mathbb{R}^{+}$ be such that $c_{j}\geq C$ for all $j\in\mathbb{Z}$ and some $C>0$. For each $j\in\mathbb{Z}$ let $\alpha_{j}$ be the translation on $\mathbb{R}$ given by $\alpha_{j}(t)=t-c_{j}$ for all $t\in\mathbb{R}$. If there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}\leq 1-\varepsilon,w_{j}\chi_{(-\infty,0)}=1$ for all $j\geq 0$ and $w_{j}=1$ for all $j<0$; then by the same arguments as in [27, Example 3.3] one can show that the conditions of Proposition 4.6 are satisfied when $W_{j}(t)=w_{j}\cdot(f\circ\alpha_{j})$ for all $f\in L^{2}(\mathbb{R}),j\in\mathbb{Z}$ and $H_{0}^{(j)}=H_{1}^{(j)}=C_{c}(\mathbb{R})$ for all $j\in\mathbb{Z}.$ However, by the construction of the sequence $\{w_{j}\}$, it follows that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}.$ On the other hand, if there exists an $\varepsilon>0$ such that $w_{j}\chi_{[0,\infty)}=1,w_{j}\chi_{(-\infty,0)}\geq 1+\varepsilon$ for all $j<0$ and $w_{j}=1$ for all $j\geq 0,$ then again the conditions of Proposition 4.6 will be satisfied, however, in this case we would again have by the construction of the sequence $\{w_{j}\}$ that $\|W_{j}\|,\|W_{j}^{-1}\|\geq 1$ for all $j\in\mathbb{Z}.$
@@
for all $\{x_{i}\}_{i\in\mathbb{Z}}\in\ell_{2}(\mathcal{C}).$ By [33, Proposition 1.2.4 (i)], the operator $R_{s,n}$ is a bounded linear operator on $\ell_{2}(\mathcal{C})$ for each $s\in\{1,\dots,N\}$ and $n\in\mathbb{N}.$ Let $\{T_{n,1}\}_{n\in\mathbb{N}},\dots,\{T_{n,N}\}_{n\in\mathbb{N}}$ be the families of operators on $\ell_{2}(\mathcal{C})$ as defined in the beginning of this section. For each $s\in\{1,\dots,N\}$ and $n\in\mathbb{N},$ we let $T_{n,s}\circ R_{s,n}$ be the composition of $T_{n,s}$ and $R_{s,n},$ that is
@@
Let $\mathcal{O},V_{1},\dots,V_{N}$ be non-empty open subsets of $\ell_{2}(\mathcal{C})$. Then there exist some $J,m_{1}\in\mathbb{N}$ and some $x\in\mathcal{O}$, $y^{(1)}\in V_{1},\dots,y^{(N)}\in V_{N}$ such that $\tilde{p}_{J,m_{1}}x=x$ and $\tilde{p}_{J,m_{1}}y^{(\ell)}=y^{(\ell)}$ for each $\ell\in\{1,\dots,N\}.$ Since $\{P_{m}\}_{m\in\mathbb{N}}$ forms also a right approximate unit for $\mathcal{C}$ by [33, Proposition 2.2.1], it follows that we can find some $m_{2}\in\mathbb{N}$ such that $\tilde{p}_{J,m_{1}}x\tilde{p}_{J,m_{2}}\in\mathcal{O}$ and $\tilde{p}_{J,m_{1}}y^{(\ell)}\tilde{p}_{J,m_{2}}\in V_{l}$ for each $\ell\in\{1,\dots,N\}.$ Set $m:=\text{max}\{m_{1},m_{2}\}.$ Then $\tilde{p}_{J,m}x\tilde{p}_{J,m}\in\mathcal{O}$ and $\tilde{p}_{J,m}y^{(\ell)}\tilde{p}_{J,m}\in V_{l}$ for each $\ell\in\{1,\dots,N\}.$ Put $C_{y}=\sum_{s=1}^{N}\|y^{(s)}\|$ and